How many grams of CuCO3(s) would be formed from the copper ions in 47.7 mL of 0.696 M CuSO4?
1
Expert's answer
2019-03-28T04:12:49-0400

net ionic equation Cu2+(aq) + CO32-(aq) = CuCO3(s)

  1.  Number of moles of Cu2+(aq) = CuSO4 = n =c x V = 0.696 moles/L x 0.0477 L = 0.0332 moles
  2. According to reaction stoichiometry number of moles of CuCO3 = CuSO4 = 0.0332 moles
  3. Mass of CuCO3 = MW x n = 123.55 g/mol x 0.0332 moles = 4.1 g